Use a volumetric pipet to deliver 2 mL of 4 M sodium hydroxide solution to each of seven 100-mL volumetric flasks. Add 2,4, 6, 8, and 10 mL, respectively, of the salicylic acid stock solution to 5 of the labeled flasks. Calculate the concentrations of the standard solutions. Use a pipet to place 10 mL of the unknown solution into another volumetric flask and label Unknown 1. Dilute each solution to the mark with DI water. The seventh flask that contains only sodium hydroxide and water will serve as the blank. .
Answer #1

(i) the volume of sodium hydroxide taken = 2mL   = V1 concentration = 4M = M1

final volume = V2   = 100mL   concentration = M2

Therefore

M1V1 = M2V2

M2 = 2 X 4 / 100= 0.08 M

2) the volume of sodium hydroxide taken = 4mL   = V1 concentration = 4M = M1

final volume = V2   = 100mL   concentration = M2

Therefore

M1V1 = M2V2

M2 = 4 X 4 / 100= 0.16 M

3) the volume of sodium hydroxide taken = 6mL   = V1 concentration = 4M = M1

final volume = V2   = 100mL   concentration = M2

Therefore

M1V1 = M2V2

M2 = 6 X 4 / 100= 0.24 M

4) the volume of sodium hydroxide taken = 8mL   = V1 concentration = 4M = M1

final volume = V2   = 100mL   concentration = M2

Therefore

M1V1 = M2V2

M2 = 8 X 4 / 100= 0.32 M

5) the volume of sodium hydroxide taken = 10mL   = V1 concentration = 4M = M1

final volume = V2   = 100mL   concentration = M2

Therefore

M1V1 = M2V2

M2 = 10 X 4 / 100= 0.4 M